# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
$\leq\left(2y_{1}+\frac{\delta_{y}}{3}\right)\frac{2^{L-1}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*{L-1}}_{\min}}\|\nabla G(\bm{W})\|_{F},$ (167)
where the second inequality uses (165) and (166). Next, we write $\bar{\bm{U}}_{L}$ and $\bm{I}_{d_{L}}$ as follows:
@@
where $\bar{\bm{U}}_{L}^{(i)},\bm{E}^{(i)}\in\mathbb{R}^{h_{i}\times d_{L}}$ for all $i\in[p_{Y}]$, and $\bar{\bm{U}}_{L}^{(p_{Y}+1)},\bm{E}^{(p_{Y}+1)}\in\mathbb{R}^{(d_{L}-r_{Y})\times d_{L}}$. Applying the Davis-Kahan Theorem (see Lemma 10) to the matrices $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ (considered as $\bm{A}$), $\bar{\bm{Y}}\bar{\bm{Y}}^{T}$ (considered as $\hat{\bm{A}}$), $\bar{\bm{U}}_{L}^{(i)^{T}}$ (considered as $\bm{V}$) , and $\bm{E}^{(i)^{T}}$ (considered as $\hat{\bm{V}}$) for each $i\in[p_{Y}+1]$, there exist orthogonal matrices $\hat{\bm{U}}_{L}^{(i)}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$ and $\hat{\bm{U}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{L}-r_{Y}}$ such that
where $\lambda_{i}$ is the $i$-th largest non-repeated eigenvalue of $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ and $\lambda_{0}=\infty$ and $\lambda_{p_{Y}+2}=-\infty$ by convention. Here, according to the SVD of $Y$ in (22), we compute $\lambda_{i}=y_{s_{i}}^{2}$ for $i\in[p_{Y}]$ and $\lambda_{p_{Y}+1}=0$ for $i\in[p_{Y}]$. Using (167) and the following inequalities
(168) yields
@@
$\left\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\right\|_{F}\leq\frac{2^{L+1}(6y_{1}+\delta_{y})\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}}\|\nabla G(\bm{W})\|_{F}.$
Therefore, we have
$\left\|\bar{\bm{U}}_{L}-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\right\|_{F}$ $\leq\sum_{i=1}^{p_{Y}+1}\left\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F},$ (169)
@@
$\eta_{5}:=\frac{2^{L+1}(6y_{1}+\delta_{y})(p_{Y}+1)\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}}.$ (170)
According to definition of $\tilde{\bm{U}}_{L}$ in (85), we have
$\left\|\bm{U}_{L}-\tilde{\bm{U}}_{L}\right\|_{F}=\left\|\bar{\bm{U}}_{L}-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F},$
where the second equality uses the definition of $\bar{\bm{U}}_{L}$ in (4.2) and the inequality is due to (169). Using the same argument in (167)-(169) to $\bar{\bm{V}}_{1}^{T}\bm{Y}^{T}\bm{Y}\bar{\bm{V}}_{1}$ and $\bar{\bm{Y}}^{T}\bar{\bm{Y}}$, we conclude that there exist orthogonal matrices $\hat{\bm{V}}_{1}^{(i)}\in\mathcal{O}^{h_{i}}$ for each $i\in[p]$ and $\hat{\bm{V}}_{1}^{(p+1)}\in\mathcal{O}^{d_{0}-r_{Y}}$ such that
$\left\|\bar{\bm{V}}_{1}-\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F}.$ (171)
Moreover, we have
$\ \left\|\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\bm{Y}\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}-\bm{Y}\right\|_{F}$
$\leq$ $\ \left\|\left(\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)-\bar{\bm{U}}_{L}\right)\bm{Y}\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}\right\|_{F}+$
$\ \left\|\bar{\bm{U}}_{L}\bm{Y}\left(\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)-\bar{\bm{V}}_{1}\right)^{T}\right\|_{F}+\|\bar{\bm{U}}_{L}\bm{Y}\bar{\bm{V}}_{1}^{T}-\bar{\bm{Y}}\|_{F}+\|\bar{\bm{Y}}-\bm{Y}\|_{F}$
@@
$\leq$ $\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*L-1}_{\min}}+2y_{1}\eta_{5}\right)\|\nabla G(\bm{W})\|_{F},$
where the last inequality uses (165), (169), and (171). This, together with the block structure of $\bm{Y}$ in (22) and (23), yields
@@
$\ \left\|\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)-\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)\right\|_{F}$
$=$ $\ \frac{1}{y_{s_{p_{Y}}}}\left\|\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\bm{Y}\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}-\bm{Y}\right\|_{F}$
Using the definition of $\tilde{\bm{V}}_{1}$, we compute
$\|\bm{V}_{1}-\tilde{\bm{V}}_{1}\|_{F}$ $=\left\|\bar{\bm{V}}_{1}-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}$
$\leq\left\|\bar{\bm{V}}_{1}-\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}+$
$\quad\ \left\|\mathrm{BlkDiag}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)\right\|_{F}$
$\leq\left(\eta_{5}+\frac{1}{y_{s_{p_{Y}}}}\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*L-1}_{\min}}+2y_{1}\eta_{5}\right)\right)\|\nabla G(\bm{W})\|_{F},$
where the last inequality follows from (171) and the above inequality.
The rest of the proof is devoted to showing that $\bm{\Pi}$ satisfies $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\cal B$. According to (166) and the definition of $\bar{\bm{Y}}$ in (4.2), we have $\left\|\bm{c}-\bm{\Pi}^{T}\bm{y}\right\|\leq{\delta_{y}}/{3},$ where $\bm{y}:=(y_{1},y_{2},\dots,y_{d_{\min}})$. This implies that there exists a permutation $\pi:[d_{\min}]\to[d_{\min}]$ such that [^1] Let $\pi:[d_{\min}]\rightarrow[d_{\min}]$ denote a permutation of the elements in $[d_{\min}]$. Note that there is an one-to-one correspondence between permutation matrix $\bm{\Pi}\in\mathbb{R}^{d_{\min}\times d_{\min}}$ and a permutation $\pi$, i.e., $\Pi_{ij}=1$ if $j=\pi(i)$ and $\Pi_{ij}=0$ otherwise for each $i\in[d_{\min}]$. Now, suppose that $\bm{\Pi}$ corresponds to $\pi$ and $\bm{\Pi}^{T}$ corresponds to $\pi^{-1}$.
$\left|c_{i}-y_{\pi^{-1}(i)}\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[d_{\min}].$ (172)
Noting that $\varphi(x)$ is continuous and differentiable at $x\neq 0$ and $\delta^{*}>0$, there exists a positive constant $\delta_{1}>0$ such that for all $|x-\sigma_{i}^{*}|\leq\min\{\delta_{1},\frac{\delta_{\sigma}}{3}\}$ and all $i\in[r_{\sigma}]$,
$\left|\varphi(x)-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3}.$ (173)
Using Weyl’s inequality, (83a), and (77a) (resp. 79), we have
$\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\{\delta_{1},\frac{\delta_{\sigma}}{3}\},$

### v3

@@ 3.2.2 Proof of the Error Bound
we have from (105) that $\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\|_{F}\leq\frac{2^{L+1}(6y_{1}+\delta_{y})\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ Therefore, we have
$\left\|\bar{\bm{U}}_{L}\!-\!\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\right\|_{F}$ $\!\leq\!\sum_{i=1}^{p_{Y}+1}\left\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F}.$ (106)
According to definition of $\tilde{\bm{U}}_{L}$ in (91), we have
@@
Next we prove the second inequality in (89). Using the same argument in (104)-(106) to $\bar{\bm{V}}_{1}^{T}\bm{Y}^{T}\bm{Y}\bar{\bm{V}}_{1}$ and $\bar{\bm{Y}}^{T}\bar{\bm{Y}}$, we conclude that there exist orthogonal matrices $\hat{\bm{V}}_{1}^{(i)}\in\mathcal{O}^{h_{i}}$ for all $i\in[p]$ and $\hat{\bm{V}}_{1}^{(p+1)}\in\mathcal{O}^{d_{0}-r_{Y}}$ such that
$\left\|\bar{\bm{V}}_{1}-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F}.$ (107)
Using the block structure of $\bm{Y}$ in (4), we have
$\ \left\|\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)\right\|_{F}$
$=$ $\ \frac{1}{y_{s_{p_{Y}}}}\left\|\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\bm{Y}\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}-\bm{Y}\right\|_{F}$
$\leq$ $\ \frac{1}{y_{s_{p_{Y}}}}\left(\left\|\left(\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)-\bar{\bm{U}}_{L}\right)\bm{Y}\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}\right\|_{F}\right$.
$\hskip 18.49988pt~~~\left.+\left\|\bar{\bm{U}}_{L}\bm{Y}\left(\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)-\bar{\bm{V}}_{1}\right)^{T}\right\|_{F}+\|\bar{\bm{U}}_{L}\bm{Y}\bar{\bm{V}}_{1}^{T}-\bar{\bm{Y}}\|_{F}+\|\bar{\bm{Y}}-\bm{Y}\|_{F}\right)$
where the last inequality is due to (102), (106), and (107). Using the definition of $\tilde{\bm{V}}_{1}$, we compute
$~~~~\|\bm{V}_{1}-\tilde{\bm{V}}_{1}\|_{F}=\left\|\bar{\bm{V}}_{1}-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}$
$\leq\left\|\bar{\bm{V}}_{1}-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}$
$\quad+\left\|\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)\right\|_{F}$
[…]
That is, the second inequality in (89) holds.
The rest of the proof is devoted to showing that $\bm{\Pi}$ satisfies $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\cal$ B. According to (103) and the definition of $\bar{\bm{Y}}$ in (101), we have $\left\|\bm{\beta}-\bm{\Pi}^{T}\bm{y}\right\|\leq{\delta_{y}}/{3},$ where $\bm{y}:=(y_{1},y_{2},\dots,y_{d_{\min}})$. This implies that there exists a permutation $\pi:[d_{\min}]\to[d_{\min}]$ such that
$\left|\beta_{i}-y_{\pi^{-1}(i)}\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[d_{\min}].$ (108)
Using Weyl’s inequality and (86), we have
@@
$\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\delta_{1},\frac{\delta_{\sigma}}{3}\right\},$
This, together with (100) and Lemma B.5 in Appendix B.3, implies
$\left|\beta_{i}-\varphi(\sigma_{i}^{*})\right|=\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[r_{\sigma}].$ (109)
According to (26), there exists $k_{i}\in[d_{\min}]$ such that $\varphi(\sigma_{i}^{*})=y_{k_{i}}$ for all $i\in[r_{\sigma}]$. For all $i\in[r_{\sigma}]$, we have
[…]
Combining this with the definition of $\delta_{y}$ in (18) yields $\varphi(\sigma_{i}^{*})=y_{k_{i}}=y_{\pi^{-1}(i)}$ for all $i\in[r_{\sigma}]$. Thus, we have
For each $i\in\{r_{\sigma}+1,\dots,d_{\min}\}$, we have $\sigma^{*}_{i}=0$, and it is trivial to see that the above equation holds. Thus, we have $(\sigma_{\pi(i)}^{*})^{2L-1}-\sqrt{\lambda}\,y_{i}\,(\sigma_{\pi(i)}^{*})^{L-1}+\lambda\,\sigma^{*}_{\pi(i)}=0,\ \sigma_{\pi(i)}^{*}\geq 0,\ \forall i\in[d_{\min}].$ This, together with the definition of $\mathcal{A}$ in (26), yields $\bm{a}:=\bm{\Pi}\bm{\sigma}^{*}=\left(\sigma_{\pi(1)}^{*},\ldots,\sigma_{\pi(d_{\min})}^{*}\right)\in\mathcal{A}$. It follows from $\bm{\sigma}^{*}=\bm{\Pi}^{T}\bm{a}$ and the definition of $\mathcal{B}$ in (27) that $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\mathcal{B}$.
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|$ $=\left|\beta_{i}-y_{\pi^{-1}(i)}\right|\leq\|\bm{\beta}-\bm{\Pi}^{T}\bm{y}\|=\|\bm{\Pi}\bm{\beta}-\bm{y}\|$
[…]
Under Assumption 2, by Lemma A.2 in Appendix A.2, we have $\varphi^{\prime}(\sigma_{i}^{*})\neq 0$ for each $i\in[r_{\sigma}]$, so that $\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|>0$. This, together with Lemma B.5 (see Appendxi B.3), yields that for all $i\in[r_{\sigma}]$,
$\left|\varphi^{\prime}(x)-\varphi^{\prime}(\sigma_{i}^{*})\right|\leq\frac{|\varphi^{\prime}(\sigma_{i}^{*})|}{2},\quad\text{if }|x-\sigma_{i}^{*}|\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\}.$ (111)
By Weyl’s inequality, we have
[…]
Applying the mean value theorem to $\varphi(\cdot)$, there exists $x$ satisfying $\sigma_{i}(\bm{W}_{L})\leq x\leq\sigma_{i}^{*}$ such that
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|=\left|\varphi^{\prime}(x)\right|\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\geq\frac{\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}{2}\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|,$
where the inequality follows from (111) and $|x-\sigma_{i}^{*}|\leq\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\delta_{2}$ due to (112). This, together with (3.2.2), yields for all $i\in[r_{\sigma}]$,
Using this and (37) in Lemma 2, we obtain the desired ∎
Now we are ready to prove the error bound for Problem (12).
Theorem 3.
Suppose that Assumptions 1 and  2 hold. Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and $\bm{W}$ be arbitrary such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\epsilon_{\bm{\sigma}^{*}}$. Then we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\kappa_{\bm{\sigma}^{*}}\|\nabla G(\bm{W})\|_{F},$ (113)
@@
$\epsilon_{\bm{\sigma}^{*}}$ $:=\begin{cases}\min\!\left\{\begin{array}[]{@{}l@{}}\frac{\delta_{\sigma}}{3},\;\delta_{1},\;\delta_{2},\;\frac{\delta_{y}\sqrt{\lambda}\,(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}L_{G}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},\\[4.30554pt] \quad\frac{\sqrt{\lambda}}{\sqrt{3(\sqrt{\lambda}+y_{1})}}\Biggl(\min\!\bigl\{\,\min_{i\in[s_{p_{Y}}]}\lvert\sqrt{\lambda}-y_{i}\rvert,\;\sqrt{\lambda}\,\bigr\}\Biggr)^{\!1/2},\\[5.16663pt] \quad\frac{\sqrt{2\lambda}\,\min\!\bigl\{\,\min_{i\in[s_{p_{Y}}]}\lvert\sqrt{\lambda}-y_{i}\rvert,\;\sqrt{\lambda}\,\bigr\}\,\sigma^{*}_{\min}}{12\,c_{2}\,L_{G}}\end{array}\right\},&\text{if}\ L=2,\\[43.05542pt] \min\!\Biggl\{\frac{\delta_{\sigma}}{3},\;\delta_{1},\;\delta_{2},\;\frac{\delta_{y}\sqrt{\lambda}\,(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}L_{G}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},\;\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{\!\frac{1}{L-2}}\Biggr\},&\text{if}\ L\geq 3,\end{cases}$ (114)
@@
$\kappa_{\bm{\sigma}^{*}}$ $:=\sqrt{L}\left(\frac{9\sigma_{\max}^{*2}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}+c_{3}\sqrt{d_{\max}-r_{\sigma}}+c_{4}\sigma_{\max}^{*}+c_{5}\sqrt{r_{\sigma}}\right).$ (115)
All the explicit expressions of the constants above can be found in Table 1.
Proof.
By the fact that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\epsilon_{\bm{\sigma}^{*}}$, together with 3, we obtain that there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48) holds. Combining this with Section B.4 in Appendix D, we have that (86) and (70) (resp., (72)) hold when $L=2$ (resp., $L\geq 3$). Moreover, it follows from 4 and 5 that (73), (89), (94) hold, where $\tilde{\bm{U}}_{L}$, $\tilde{\bm{V}}_{1}$, and $\hat{\bm{T}}$ are respectively defined in (91), (92), and (93). Recall that $\bm{U}_{l}\in\mathcal{O}^{d_{l}}$ and $\bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$ for all $l\in[L]$ are introduced in (3.2.1). For ease of exposition, we define $\bm{\Sigma}_{l}^{*}:=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma}^{*}),\bm{0})\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and
$\hat{\bm{W}}_{1}:=\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T},\qquad\hat{\bm{W}}_{l}:=\bm{U}_{l}\bm{\Sigma}_{l}^{*}\bm{V}_{l}^{T},\ l=2,\dots,L-1,\qquad\hat{\bm{W}}_{L}:=\tilde{\bm{U}}_{L}\bm{\Sigma}_{L}^{*}\bm{V}_{L}^{T}.$ (116)
Now, we compute
$\|\hat{\bm{W}}_{1}-\bm{W}_{1}\|_{F}$ $=\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{T}\right\|_{F}$
$\leq\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}^{*}\bm{V}_{1}^{T}\right\|_{F}+\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\bm{V}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{T}\right\|_{F}$
$\leq\sigma_{\max}^{*}\|\tilde{\bm{V}}_{1}-\bm{V}_{1}\|_{F}+\|\bm{\Sigma}_{1}^{*}-\bm{\Sigma}_{1}\|_{F}$
For all $l=2,\dots,L-1$, we compute
[…]
This, together with (3.2.2) and (118), yields
Next, we further define
$\bm{T}_{l}:=\mathrm{BlkD}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}),\ l=2,\dots,L,\ \bm{Q}_{L}:=\bm{V}_{L},$ (120)
$\bm{W}_{1}^{*}:=\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\tilde{\bm{V}}_{1}^{T},\ \bm{W}_{l}^{*}:=\bm{Q}_{l+1}\bm{\Sigma}_{l}^{*}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{W}_{L}^{*}:=\tilde{\bm{U}}_{L}\bm{\Sigma}_{L}^{*}\bm{Q}_{L}^{T},$ (121)
@@
$\textstyle\bm{Q}_{l}:=\bm{V}_{l}\mathrm{BlkD}\left(\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(1)}\right),\dots,\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(p)}\right),\bm{I}_{d_{l-1}-r_{\sigma}}\right),\ l=2,\dots,L-1.$ (122)
Using 1, one can verify that $\bm{W}^{*}=(\bm{W}_{1}^{*},\bm{W}_{2}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}.$ For each $l=2,\dots,L-2$, we compute
$\|\hat{\bm{W}}_{l}-\bm{W}_{l}^{*}\|_{F}$ $=\left\|\bm{U}_{l}\bm{\Sigma}_{l}^{*}\bm{V}_{l}^{T}-\bm{Q}_{l+1}\bm{\Sigma}_{l}^{*}\bm{Q}_{l}^{T}\right\|_{F}$
[…]
where the last equality uses the diagonal block forms of $\bm{T}_{l}$ and $\bm{\Sigma}_{l}^{*}$. Using (116) and (121), we compute
[…]
where the last equality follows from the froms of $\bm{T}_{l}$ in (120), $\bm{Q}_{2}$ in (122), and $\hat{\bm{T}}$ in (93). This, together with (3.2.2) and the fact that $\bm{W}_{L}^{*}=\hat{\bm{W}}_{L}$ due to $\bm{Q}_{L}=\bm{V}_{L}$, yields
$\|\hat{\bm{W}}-\bm{W}^{*}\|_{F}\leq\frac{9\sigma_{\max}^{*2}\sqrt{L-1}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ (124)